OpenStudy (anonymous):

4 over 5y-3 = 2 over 3y +7 solve

OpenStudy (sasogeek):

\[\large \frac{4}{5y-3} = \frac{2}{3y+7} \]\[\large 12y+28=10y-6 \]\[\large 2y=-34 \]\[\large y=-17 \]